OpenStudy (anonymous):

The width w of a rectangular swimming pool is 7x^2. The area A of the pool is 7x^3 – 42x^2. What is an expression for the length of the pool?

Parth (parthkohli):

\[\rm area = length \times width\]So,\[\rm length = \dfrac{area}{width}\]Can you figure everything else out?

OpenStudy (anonymous):

no but i will try lol

Parth (parthkohli):

What is the length?

OpenStudy (anonymous):

\[length=f(x)\]

Parth (parthkohli):

What is the width?

OpenStudy (anonymous):

width is 7x^2

OpenStudy (anonymous):

so plug that into the equation \[A=LW=7x^3-42x^2\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

then divide by W to get \[\frac{7x^3-42x^2}{W}\]